# Description length of canonical and microcanonical models

### v3

@@ Acknowledgements
[…]
$=-\ln\int_{0}^{1}p^{l^{*}}(1-p)^{nm-l^{*}}\frac{1}{\pi\sqrt{p(1-p)}}$
[…]
@@
$\text{DL}^{\text{Bayes-Jeffreys}}_{\text{can}}(\mathbf{G}^{*})=-\ln\frac{\Gamma\left(l^{*}+\frac{1}{2}\right)\Gamma\left(nm-l^{*}+\frac{1}{2}\right)}{\pi(nm)!},$ (S19)
@@
$\text{DL}^{\text{Bayes-Jeffreys}}_{\text{can}}(\mathbf{G}^{*})=nm\cdot h\left(\frac{l^{*}}{nm}\right)+\frac{1}{2}\ln\frac{\pi nm}{2}+O\left(1/nm\right).$ (S20)
The last expression is compared with the asymptotic expansion of the canonical NML description length $\text{DL}^{\text{NML}}_{\text{can}}$, obtained as the difference between the asymptotic expansion of (S8) and the canonical log-likelihood:
$\text{DL}^{\text{NML}}_{\text{can}}(\mathbf{G}^{*})=nm\cdot h\left(\frac{l^{*}}{nm}\right)+\frac{1}{2}\ln\frac{\pi nm}{2}+\frac{2}{3}\sqrt{\frac{2}{\pi nm}}+o(1).$ (S21)
Finally, Equation (82) is obtained as the difference between (S21) and (S20)
S2.2 Proving identities
@@
Proving Identity (85).
We begin with identity (85), namely:
$\text{DL}_{\text{can}}^{\text{NML}}=\text{DL}_{\text{mic}}^{\text{Bayes-Rissanen}},$ (S22)
@@
$\hat{W}(\mathbf{c})$ $=\frac{\sum_{\mathbf{G}:\mathbf{c}(\mathbf{G})=\mathbf{c}}P_{\text{can}}(\mathbf{G};\hat{\bm{\theta}}(\mathbf{G})}{\sum_{\mathbf{G}}P_{\text{can}}(\mathbf{G};\hat{\bm{\theta}}(\mathbf{G}))}$ (S23)
$=\frac{\Omega(\mathbf{c})\mathcal{L}_{\text{can}}(\mathbf{c})}{\sum_{\mathbf{G}}\mathcal{L}_{\text{can}}(\mathbf{G})}.$
@@
$\text{DL}_{\text{mic}}^{\text{Bayes-Rissanen}}(\mathbf{G}^{*})$ $=\ln\Omega(\mathbf{c}^{*})-\ln\hat{W}(\mathbf{c}^{*})=$ (S24)
@@
which proves identity (85).
Proving Identity (86).
We will now prove identity (86), namely:
$\text{DL}_{\text{mic}}^{\text{NML}}=\text{DL}_{\text{can}}^{\text{Bayes-Uniform}},$ (S25)
for the specific MEMs considered in this work, starting from the case of one global constraint. The uniform prior on the mean-value parameter $p$ reads
$w^{\text{Uniform}}(p)=1\quad\text{for }p\in[0,1].$ (S26)
@@
[…]
@@
$=\text{DL}_{\text{mic}}^{\text{NML}}(\mathbf{G}^{*}),$
where the integral is computed by integrating by parts $l^{*}$ times. This proves identity (86) for the case of one global constraint.
The identity holds as well for the case of $n$ one-sided local constraint, with the uniform prior reading
$w^{\text{Uniform}}(\mathbf{p})=1\quad\text{for }\mathbf{p}\in[0,1]^{n}$ (S28)
@@
[…]
@@
$=\text{DL}_{\text{mic}}^{\text{NML}}(\mathbf{G}^{*}),$ (S30)
which proves identity (86).

### v4

@@ C.1 Bayesian-Jeffreys description lengths
$\text{DL}^{\text{Bayes-Jeffreys}}_{\text{can}}(\mathbf{G}^{*})=-\ln\int_{0}^{1}P_{\text{can}}(\mathbf{G}^{*};p)J(p)dp\\ =-\ln\frac{1}{\pi}\int_{0}^{1}p^{l^{*}-\frac{1}{2}}(1-p)^{nm-l^{*}-\frac{1}{2}}dp.$ (102)
@@
$\text{DL}^{\text{Bayes-Jeffreys}}_{\text{can}}(\mathbf{G}^{*})=\\ =-\ln\frac{\Gamma\left(l^{*}+\frac{1}{2}\right)\Gamma\left(nm-l^{*}+\frac{1}{2}\right)}{\pi(nm)!},$ (105)
@@
$\text{DL}^{\text{Bayes-Jeffreys}}_{\text{can}}(\mathbf{G}^{*})=nm\cdot h\left(\frac{l^{*}}{nm}\right)\\ +\frac{1}{2}\ln\frac{\pi nm}{2}+\Theta\left(1/nm\right).$ (106)
The last expression is compared with the asymptotic expansion of the canonical NML description length $\text{DL}^{\text{NML}}_{\text{can}}$, obtained as the difference between the asymptotic expansion of (94) and the canonical log-likelihood:
$\text{DL}^{\text{NML}}_{\text{can}}(\mathbf{G}^{*})=nm\cdot h\left(\frac{l^{*}}{nm}\right)\\ +\frac{1}{2}\ln\frac{\pi nm}{2}+\frac{2}{3}\sqrt{\frac{2}{\pi nm}}+o(1).$ (107)
Finally, Equation (80) is obtained as the difference between (107) and (106)
C.2 Proving identities
@@
Proving Identity (83). We begin with identity (83), namely:
$\text{DL}_{\text{can}}^{\text{NML}}=\text{DL}_{\text{mic}}^{\text{Bayes-Rissanen}},$ (108)
@@
$\hat{W}(\mathbf{c})$ $=\frac{\sum_{\mathbf{G}:\mathbf{c}(\mathbf{G})=\mathbf{c}}P_{\text{can}}(\mathbf{G};\hat{\bm{\theta}}(\mathbf{G})}{\sum_{\mathbf{G}}P_{\text{can}}(\mathbf{G};\hat{\bm{\theta}}(\mathbf{G}))}$
$=\frac{\Omega(\mathbf{c})\mathcal{L}_{\text{can}}(\mathbf{c})}{\sum_{\mathbf{G}}\mathcal{L}_{\text{can}}(\mathbf{G})}.$ (109)
@@
$\text{DL}_{\text{mic}}^{\text{Bayes-Rissanen}}(\mathbf{G}^{*})$ $=\ln\Omega(\mathbf{c}^{*})-\ln\hat{W}(\mathbf{c}^{*})$ (110)
@@
which proves the identity.
Proving Identity (84). We will now prove identity (84), namely:
$\text{DL}_{\text{mic}}^{\text{NML}}=\text{DL}_{\text{can}}^{\text{Bayes-Uniform}},$ (111)
for the specific maximum-entropy models considered in this work, starting from the case of one global constraint. The uniform prior on the mean-value parameter $p$ reads
$w^{\text{Uniform}}(p)=1\quad\text{for }p\in[0,1].$ (112)
@@
$\text{DL}_{\text{can}}^{\text{Bayes-Uniform}}(\mathbf{G}^{*})$ $=-\ln\int_{0}^{1}P_{\text{can}}(\mathbf{G}^{*};p)dp$
$=-\ln\int_{0}^{1}p^{l^{*}}(1-p)^{nm-l^{*}}dp$
@@
$=\text{DL}_{\text{mic}}^{\text{NML}}(\mathbf{G}^{*}),$ (113)
where the integral is computed by integrating by parts $l^{*}$ times. This proves the identity for the case of one global constraint.
The identity holds as well for the case of $n$ one-sided local constraints, with the uniform prior reading
$w^{\text{Uniform}}(\mathbf{p})=1\quad\text{for }\mathbf{p}\in[0,1]^{n}.$ (114)
@@
$\text{DL}_{\text{can}}^{\text{Bayes-Uniform}}(\mathbf{G}^{*})$ $=-\ln\int_{0}^{1}P_{\text{can}}(\mathbf{G}^{*};p)dp$
$=-\sum_{i=1}^{n}\ln\int_{0}^{1}p_{i}^{r_{i}^{*}}(1-p)^{m-r_{i}^{*}}dp_{i}$
@@
$=\text{DL}_{\text{mic}}^{\text{NML}}(\mathbf{G}^{*}),$ (115)
which proves the identity for the case of $n$ one-sided local constraints.